A siren mounted on the roof of a firehouse emits sound at a frequency of 900 Hz. A steady wind is blowing with a speed of 12.0 m/s. Firefighters are approaching the siren from various directions at 15.0 m/s. What frequency does a firefighter hear if he is approaching from an upwind position so that he is moving in the direction in which the wind is blowing? (Speed of sound in calm air to be 340 m/s 1.
Answer #1

The frequency heard by the fire fighter is given by

f'=\left ( \frac{v+v_{o}}{v-v_{s}} \right )f=\left ( \frac{340+0}{340-15} \right )\times900

f'=941.5 Hz